# Cohen–Macaulay Complexes, Duality Groups, and the dualizing module of ${\rm{Out}}(F_{N})$

### v1

@@ 3.4. The dualizing module of ${\rm{Out}}(F_{N})$
The interior of the thickened spine is homeomorphic to Outer space, so $h_{*}(\mathcal{\mathcal{L}_{N}})|_{\mathcal{X}_{N}}=h^{*}(\mathcal{CV}_{N})|_{\mathcal{X}_{N}}$. As $\mathcal{L}_{N}$ is contractible, $H_{3N-4-k}(\mathcal{L}_{N},\partial\mathcal{L}_{N};\mathbb{Z})\cong H_{3N-5-k}(\partial\mathcal{L}_{N};\mathbb{Z})$ by the long exact sequence in homology. Finally $\partial\mathcal{L}_{N}$ is homotopy equivalent to $\partial\mathcal{FS}_{N}$ by Theorem 3.8. The same argument goes through for the final isomorphism, using $\mathcal{K}_{N}$ in place of $\mathcal{L}_{N}$.
@@
In dimension two, links of points are homeomorphic to graphs, and we will see that the local Cohen–Macaulay condition is equivalent to the graphs that appear as vertex links being connected and nondegenerate (not a single point). A related condition called visible irreducibility, appears in work of Louder and Wilton [LW24] (reformulated as below in [Wil22, Remark 4.2]) . Informally this says that there are no obvious local simplifications one can perform on the 2-complex. More precisely, a complex is visibly irreducible if:
@@
Otherwise, a complex is visibly reducible. We now turn to Theorem E from the introduction, which is restated below.
@@
Let $Y$ be simplicial 2-complex. Then $Y$ is a local homology CM complex if and only if it satisfies (i) and (ii). Furthermore, the following are equivalent:
@@
For the first part, we want to discuss when the local homology $h_{*}(\sigma)$ is concentrated in degree 2 at every simplex $\sigma\in Y$. This is clearly the case for 2-simplices. Let $V\in Y$ be a vertex and let $E\in Y$ be an edge containing $V$. Let $\Gamma=\mathrm{lk}_{Y}(V)$ be the vertex link. Every adjacent edge at $V$ determines a vertex $v_{E}\in\Gamma$ and 2-simplices containing $V$ give edges in $\Gamma$. Consequently, the chain complex $C_{\bullet}^{V}(Y)$ (as defined in Section 2.2) is isomorphic to an augmented version of $C_{\bullet-1}(\Gamma)$. This implies that $h_{1}(V)\cong\tilde{H}_{0}(\Gamma)$ and $h_{2}(V)\cong H_{1}(\Gamma)$. Hence the CM condition holds at each vertex if and only if each vertex link is connected. The CM condition holds at an edge $E$ of $Y$ as long as $h_{1}(E)$ vanishes, or equivalently $E$ is contained in a face. The edge $E$ is contained in a face unless its vertex $v_{E}\in\Gamma$ is isolated. A connected graph only contains an isolated vertex if it is a single point. Putting all of the above together, $Y$ is a local homology CM complex if and only if the vertex links of $Y$ satisfy (i) and (ii).
As there is no torsion in local (co)homology when the complex is locally CM, equivalence of the second and third bullet points follows from the universal coefficient theorem. Therefore we will focus on the local homology sheaf. Firstly, if $F$ is a 2–simplex in $Y$ and $E$ is an edge of $F$, the map $h_{2}(E)\to h_{2}(F)$ is surjective if and only if the local homology $h_{2}(E)$ does not vanish. This happens if and only if $E$ is contained in at least two 2–simplices. This in turn is equivalent to vertex links not containing leaves.

### v2

@@ 3.4. The dualizing module of ${\rm{Out}}(F_{N})$
where $h_{*}(\mathcal{CV}_{N})|_{\mathcal{X}_{N}}$ and and $h_{*}(\mathcal{CV}^{r}_{N})|_{{\mathcal{X}_{N}^{r}}}$ are the restrictions of the local homology sheaf of Outer space (and reduced Outer space) to the spine (and its reduced version, respectively).
@@
The interior of the thickened spine is homeomorphic to Outer space, so $h_{*}(\mathcal{\mathcal{L}_{N}})|_{\mathcal{X}_{N}}=h_{*}(\mathcal{CV}_{N})|_{\mathcal{X}_{N}}$. As $\mathcal{L}_{N}$ is contractible, $H_{3N-4-k}(\mathcal{L}_{N},\partial\mathcal{L}_{N};\mathbb{Z})\cong H_{3N-5-k}(\partial\mathcal{L}_{N};\mathbb{Z})$ by the long exact sequence in homology. Finally $\partial\mathcal{L}_{N}$ is homotopy equivalent to $\partial\mathcal{FS}_{N}$ by Theorem 3.8. The same argument goes through for the final isomorphism, using $\mathcal{K}_{N}$ in place of $\mathcal{L}_{N}$. ∎
@@
4. Bounding the top-dimensional cohomology of ${\rm{Out}}(F_{N})$
In this section, we see how cosheaf homology gives a resolution for the dualizing module of ${\rm{Out}}(F_{N})$, and use this to prove the following:
Theorem 4.1.
Let $\rho\in{\mathcal{X}_{N}^{r}}$ be a rose in the spine of reduced outer space, and let ${\rm{Stab}}(\rho)<{\rm{Out}}(F_{N})$ be the stabilizer of this rose. Then
$\dim(H^{2N-3}({\rm{Out}}(F_{N});\mathbb{Q}))\leq\dim((h^{2N-3}(\rho)\otimes\mathbb{Q})_{{\rm{Stab}}(\rho)}),$
where $h^{2N-3}(\rho)$ is the local cohomology of the reduced spine at this rose.
Recall that a rose is a point $T$ in the spine such that $T/F_{N}$ is a graph with one vertex and $N$ loops, and we use $M_{G}$ to denote the group of coinvariants of a $G$-module $M$. As ${\rm{Out}}(F_{N})$ is a virtual duality group, it is a duality group of dimension $32N-$ 3 over $\mathbb{Q}$ [Bie81, Theorem 9.9]. In particular this implies:
$H^{2N-3}({\rm{Out}}(F_{N});\mathbb{Q})\cong H_{0}({\rm{Out}}(F_{N});D\otimes\mathbb{Q})=(D\otimes\mathbb{Q})_{{\rm{Out}}(F_{N})},$
where $D$ is the dualizing module of ${\rm{Out}}(F_{N})$. To prove Theorem 4.1, we will show that with integer coefficients one has
(7) $D_{{\rm{Out}}(F_{N})}\subseteq h^{2N-3}(\rho)_{{\rm{Stab}}(\rho)},$
which implies the dimensional inequality after passing to rational coefficients. As above, the dualizing module of ${\rm{Out}}(F_{N})$ is
$D=H^{2N-3}_{c}({\mathcal{X}_{N}^{r}};\mathbb{Z}).$
As we are in the top dimension, the coboundary map is trivial so $H^{2N-3}_{c}({\mathcal{X}_{N}^{r}};\mathbb{Z})$ is generated by cohomology classes of the form $[\alpha^{*}]$, as $\alpha$ varies over the top-dimensional simplices of ${\mathcal{X}_{N}^{r}}$. Viewing $G={\rm{Out}}(F_{N})$ as acting on ${\mathcal{X}_{N}^{r}}$ on the left, this gives $H^{2N-3}_{c}({\mathcal{X}_{N}^{r}};\mathbb{Z})$ the usual right $G$-module structure by
$[\alpha^{*}]g=[g^{-1}(\alpha)^{*}].$
By [Vog90], the spine is locally Cohen–Macaulay, so we can apply CM duality (Theorem 2.1) to obtain an isomorphism
$[{\mathcal{X}_{N}^{r}}]\cap-\colon H^{k}_{c}({\mathcal{X}_{N}^{r}};\mathbb{Z})\xrightarrow{\cong}H_{2N-3-k}({\mathcal{X}_{N}^{r}};h^{2N-3}),$
where $h^{2N-3}$ is the local cohomology cosheaf of the spine and $[{\mathcal{X}_{N}^{r}}]\in H^{\text{lf}}_{2N-3}({\mathcal{X}_{N}^{r}};h^{2N-3})$ is the fundamental class. As $H^{*}_{c}({\mathcal{X}_{N}^{r}};\mathbb{Z})$ is concentrated in degree $32N-$ 3, $H_{*}({\mathcal{X}_{N}^{r}};h^{2N-3})$ is concentrated in degree zero, so its chain complex is acyclic. This gives a resolution
(8) $\cdots\to C_{1}({\mathcal{X}_{N}^{r}};h^{2N-3})\to C_{0}({\mathcal{X}_{N}^{r}};h^{2N-3})\to D\to 0$
of the dualizing module. The modules in this chain complex are (see Equation (1))
$C_{k}({\mathcal{X}_{N}^{r}};h^{2N-3})=\oplus_{\sigma\in({\mathcal{X}_{N}^{r}})_{k}}h^{2N-3}(\sigma),$
where $h^{2N-3}(\sigma)=H^{2N-3}({\mathcal{X}_{N}^{r}},{\mathcal{X}_{N}^{r}}-\mathring{\mathrm{st}}(\sigma))$ is the local homology at $\sigma$. To distinguish where local homology classes live, for $\phi\in h^{2N-3}(\sigma)$ we use $\sigma\triangleleft\phi$ to denote the corresponding element of $C_{n}({\mathcal{X}_{N}^{r}};h^{2N-3})$ in the $h^{2N-3}(\sigma)$-summand.
At the chain level, the cap product with the fundamental class $[{\mathcal{X}_{N}^{r}}]$ given in [WW23] is the $G$-equivariant (see [WW23, Section 5]) map
$c_{2N-3}\colon C^{2N-3}({\mathcal{X}_{N}^{r}};\mathbb{Z})\to C_{0}({\mathcal{X}_{N}^{r}};h^{2N-3})$
given by
(9) $c_{2N-3}(\alpha^{*})=\alpha_{0}\triangleleft[\alpha^{*}],$
where $\alpha_{0}$ is the first vertex of $\alpha$ (using the ordering/orientation on ${\mathcal{X}_{N}^{r}}$ induced by the poset structure), and $\alpha_{0}\triangleleft[\alpha^{*}]$ lives in the local cohomology at the vertex $\alpha_{0}$. The quickest way to see this is via Equation (14) on page 13 of [WW23], noting that the fundamental class $[{\mathcal{X}_{N}^{r}}]$ is represented by the infinite sum $\sum{\alpha\triangleleft[\alpha^{*}]}$, as $\alpha$ ranges over all maximal simplices. In the poset structure we use on the spine every maximal simplex $\alpha$ is given by a chain
$T_{0}<T_{1}<\cdots<T_{2N-3},$
where the initial vertex $T_{0}$ of $\alpha$ is a rose. We can leverage this to restrict the image of $c_{2N-3}$:
Proposition 4.2.
Let $R$ be the submodule of $C_{0}({\mathcal{X}_{N}^{r}};h^{2N-3})$ spanned by the local cohomology groups of the roses in ${\mathcal{X}_{N}^{r}}$. Every element of $D=C_{0}({\mathcal{X}_{N}^{r}};\mathbb{Z})/\operatorname{im}{d_{1}}$ has a representative in $R$
Proof.
The cap product $c_{2N-3}$ induces an isomorphism between $H^{2N-3}_{c}({\mathcal{X}_{N}^{r}};\mathbb{Z})$ and $H_{0}({\mathcal{X}_{N}^{r}};h^{2N-3})$. The image of $c_{2N-3}$ is contained in $R$ by (9), as the first vertex $\alpha_{0}$ of any maximal simplex $\alpha$ is a rose. In particular, $H_{0}({\mathcal{X}_{N}^{r}};h^{2N-3})$ is generated by the classes in $R$. ∎
Corollary 4.3.
Let $R$ be the submodule of $C_{0}({\mathcal{X}_{N}^{r}};h^{2N-3})$ spanned by the local cohomology groups of the roses in ${\mathcal{X}_{N}^{r}}$, and let
$d_{1}:C_{1}({\mathcal{X}_{N}^{r}};h^{2N-3})\to C_{0}({\mathcal{X}_{N}^{r}};h^{2N-3})$
be the boundary map in cosheaf homology. Then
$D\cong R/(\operatorname{im}d_{1}\cap R).$
Proof.
By CM dualiy,
$D\cong H_{0}({\mathcal{X}_{N}^{r}};h^{2N-3})\cong C_{0}({\mathcal{X}_{N}^{r}};h^{2N-3})/\operatorname{im}d_{1},$
and the inclusion $R\subset C_{0}({\mathcal{X}_{N}^{r}};h^{2N-3})$ induces an injection
$R/(\operatorname{im}d_{1}\cap R)\hookrightarrow C_{0}({\mathcal{X}_{N}^{r}};h^{2N-3})/\operatorname{im}d_{1}\cong D.$
By Proposition 4.2, every element of $D$ has a representative in $R$, so this map is also surjective. ∎
Proposition 4.4.
Let $R$ be as above and let $I$ be the augmentation ideal of the group ring of ${\rm{Out}}(F_{N})$. If $\rho$ is a rose in ${\mathcal{X}_{N}^{r}}$ and ${\rm{Stab}}(\rho)$ is its stabilizer in ${\rm{Out}}(F_{N})$ then
$R/RI=R_{{\rm{Out}}(F_{N})}\cong h^{2N-3}(\rho)_{{\rm{Stab}}(\rho)}.$
Proof.
As $R$ decomposes as a direct summand over the $h^{2N-3}(\rho)$ and ${\rm{Out}}(F_{N})$ acts transitively on the summands, it follows that
$R=\textnormal{Ind}^{{\rm{Out}}(F_{N})}_{{\rm{Stab}}(\rho)}h^{2N-3}(\rho),$
i.e. $R$ is the induced representation of the ${\rm{Stab}}(\rho)$ representation $h^{2N-3}(\rho)$ (see e.g. Proposition III.5.3 in [Bro82]). Then
$H_{*}({\rm{Out}}(F_{N});R)\cong H_{*}({\rm{Stab}}(\rho);h^{2N-3}(\rho))$
by Shapiro’s lemma [Bro82, Proposition III.6.2]. The result follows by taking =0*=0∗ = 0. ∎
Proof of Theorem 4.1.
By Corollary 4.3, we have
$D\cong\frac{R}{\operatorname{im}d_{1}\cap R},$
therefore the coinvariants of this module are given by
$D_{{\rm{Out}}(F_{N})}\cong\frac{R}{\operatorname{im}d_{1}\cap R+RI}\subset\frac{R}{RI}\cong h^{2N-3}(\rho)_{{\rm{Stab}}(\rho)}$
Theorem 4.1 is then a consequence of ${\rm{Out}}(F_{N})$ being a duality group over $\mathbb{Q}$ with rational dualizing module $D\otimes\mathbb{Q}$. ∎
Remark 4.5.
We finish this section with some remarks.
  • To show that the inclusion from Equation (7) is actually an isomorphism, it follows from the proof that it is enough to show that $RI$ contains $\operatorname{im}d_{1}\cap R$. This is definitely false in the unreduced case, and seems unlikely to be true even in the reduced setting.
  • The representation theory of the ${\rm{Stab}}(\rho)$ action on the cohomology of the link (or equivalently, the local cohomology) is well understood in terms of tree representations of the symmetric group. See [Bar90] and [RW96]. For instance, in the unreduced setting the dimension of $h^{2N-3}(\rho)$ is $(2N-2)!$ However, the structure of the link is much less well-understood in the reduced spine.
  • While we suspect it may be possible to prove Theorem 4.1 without using CM duality, the theorem was only formulated after a careful consideration of the the resolution of the dualizing module obtained in Equation (8). Furthermore, we found that the cap product in CM duality facilitated a particularly clean proof of Proposition 4.2.
@@ 4. Bounding the top-dimensional cohomology of ${\rm{Out}}(F_{N})$
In dimension two, links of points are homeomorphic to graphs, and we will see that the local Cohen–Macaulay condition is equivalent to the graphs that appear as vertex links being connected and nondegenerate (not a single point). A related condition called visible irreducibility, appears in work of Louder and Wilton [LW24] (reformulated as below in [Wil24, Remark 4.2]) . Informally this says that there are no obvious local simplifications one can perform on the 2-complex. More precisely, a complex is visibly irreducible if:
@@
Otherwise, a complex is visibly reducible. We now turn to Theorem F from the introduction, which is restated below.
@@
Let $Y$ be a simplicial 2-complex. Then $Y$ is a local homology CM complex if and only if it satisfies (i) and (ii). Furthermore, the following are equivalent:
@@
We first show that if $Y$ is locally CM then it satisfies (i) and (ii). If $Y$ is locally CM then it is pure (all of its maximal simplices are 2-simplices), so in particular every vertex is contained in a 2-simplex and (i) holds. To see that (ii) also holds, let $V\in Y$ be a vertex, and let $\mathrm{lk}_{Y}(V)$ be the vertex link. By comparing the chain complexes $C_{\bullet}^{V}(Y)$ (as defined in Section 2.2) and $\tilde{C}_{\bullet-1}(\mathrm{lk}_{Y}(V))$ (that computes the reduced homology of the link) we see that $h_{k}(V)\cong\tilde{H}_{k-1}(\mathrm{lk}_{Y}(V))$: every adjacent edge $E$ at $V$ determines a vertex $v_{E}\in\mathrm{lk}_{Y}(V)$ and 2-simplices containing $V$ give edges in $\mathrm{lk}_{Y}(V)$. If (ii) does not hold, there exists a vertex $V$ such that $\mathrm{lk}_{Y}(V)$ is a disconnected graph, so that $h_{1}(V)=\tilde{H}_{0}(\mathrm{lk}_{Y}(V))$ is nontrivial. So $Y$ being locally CM also implies (ii).
Now assume that $Y$ satisfies (i) and (ii). We want to to show that the local homology $h_{*}(\sigma)$ is concentrated in degree 2 for every simplex $\sigma\in Y$. We start with the vertices: by (i) and (ii) each vertex link is a non-empty connected graph with at least one edge, so $\tilde{H}_{*}(\mathrm{lk}_{Y}(V))$ is concentrated in degree 1, and therefore $h_{*}(V)\cong\tilde{H}_{*-1}(\mathrm{lk}_{Y}(V))$ is concentrated in degree 2. Moving on to the edges, $h_{*}(E)$ is concentrated in degree two if and only if $E$ is contained in a face. If this is not the case, $E$ will be an isolated vertex in each of its vertex links, contradicting (ii) (as there is at least one edge in each vertex link by (i)). The local homology at each 2-simplex $\sigma$ satisfies $h_{2}(\sigma)=\mathbb{Z}$ and $h_{k}(\sigma)=0$ for $k\neq 2$ by maximality. Putting all of the above together, $Y$ is a local homology CM complex if and only if the vertex links of $Y$ satisfy (i) and (ii).
We now move on the equivalence of the bullet points. As there is no torsion in local (co)homology when the complex is locally CM, equivalence of the second and third bullet points follows from the universal coefficient theorem. Therefore we will focus on the local homology sheaf. Firstly, if $F$ is a 2–simplex in $Y$ and $E$ is an edge of $F$, the map $h_{2}(E)\to h_{2}(F)$ is surjective if and only if the local homology $h_{2}(E)$ does not vanish. This happens if and only if $E$ is contained in at least two 2–simplices. This in turn is equivalent to vertex links not containing leaves.